OpenStudy (anonymous):

If f(x) = 5x + 1. what is f(-2) -11 9 -1/5 -1

OpenStudy (dinamix):

f(-2) = 5*(-2)+1 f(-2)=-10+1 f(-2)=-9

OpenStudy (dinamix):

@study_123

OpenStudy (anonymous):

Thx